the second virial coefficeint for Ar gas at 600K is 11.9 cm^3 mol^-1. Calculate the following,
molar volume of Ar gas at 600K and at 1.2 atm pressure
1
Expert's answer
2020-09-24T13:50:08-0400

Ideal:

PV = RT

V = RT/P

V = (8.314 x 600) / (1.2 x 101325)

= 0.041 m3 = 41 L/mol

Real = z x Ideal

z = 1 + B(T) / V

z = 1 + 11.9 / 41

= 1.29

Real = 1.29 x 41 = 52.9 L/mol